# Origin-Destination Pattern Effects on Large-Scale Mixed Traffic Control via Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Traffic congestion remains a major challenge for modern urban transportation, diminishing both efficiency and quality of life. While autonomous driving technologies and reinforcement learning (RL) have shown promise for improving traffic control, most prior work has focused on small-scale networks or isolated intersections. Large-scale mixed traffic control, involving both human-driven and robotic vehicles, remains underexplored. In this study, we propose a decentralized multi-agent reinforcement learning framework for managing large-scale mixed traffic networks, where intersections are controlled either by traditional traffic signals or by robotic vehicles. We evaluate our approach on a real-world network of 14 intersections in Colorado Springs, Colorado, USA, using average vehicle waiting time as the primary measure of traffic efficiency. Results demonstrate that strategically adjusting major origin-destination (OD) flow patterns can effectively reduce congestion, offering a new pathway for enhancing urban mobility.
@@ I Introduction
Traffic control has advanced considerably in recent years, with reinforcement learning (RL) emerging as a powerful tool in complex traffic scenarios. RL-based vehicle control strategies often outperform traditional traffic signal methods and have recently been extended to large-scale mixed traffic environments, especially for urban intersection management [5, 6]. Among these, multi-agent reinforcement learning (MARL) stands out for its strong performance and novel approach, attracting growing research interest [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
We conduct experiments on a large-scale traffic network of 14 intersections using the Simulation of Urban MObility (SUMO)[18]. In this network, some intersections are managed by traffic signals, while others are controlled by RVs equipped with RL capabilities. Our performance evaluation metric is average vehicle waiting time, which measures traffic efficiency by averaging the total time vehicles remain motionless near an intersection. We conduct experiments under various origin-destination (OD) flow configurations and RV penetration rates. The results indicate that variations in OD configurations, when combined with appropriate RV penetration rates, can lead to significant differences in congestion levels. These findings suggest that adjusting the primary OD flow directions could serve as a viable strategy for alleviating urban traffic congestion. Such an approach opens up new avenues for dynamic traffic management and offers valuable insights for the future development of intelligent transportation systems. The code of our work is available at https://github.com/cgchrfchscyrh/MixedTrafficControl_IROS.
@@ II Related Work
Previous studies on unsignalized intersection control have mostly focused on connected and autonomous vehicles (CAVs). Early work introduced a multi-agent system where CAVs reserve space-time slots on a first-come, first-served (FCFS) basis [23]. Later research extended FCFS to CAV platoons [24] and proposed dynamic gap adjustments to reduce conflicts [25]. Decentralized approaches, like consensus-based trajectory planning [26] and energy optimization [27], have also been explored, but most assume a fully autonomous environment. In contrast, we address mixed traffic by assuming only RVs can communicate across intersections, offering a more realistic and scalable control strategy for real-world urban networks.
@@ III-A Mixed Traffic Control at Intersections Based on RL
$(S,A,T,R,O,Z,\mu_{0},\gamma),$
@@
  • $S$ is the set of all possible traffic states;
  • $A$ denotes the available actions for each vehicle;
  • $T(s^{\prime}\mid s,a)$ defines the transition probability from state $s$ to $s^{\prime}$ given action $a$;
  • $R(s,a)$ specifies the reward for taking action $a$ in state $s$;
  • $O$ is the space of partial observations available to robot vehicles (RVs);
  • $Z(o\mid s)$ gives the probability of observing $o$ when the system is in state $s$;
  • $\mu_{0}$ is the initial state distribution;
  • $\gamma$ is the discount factor for future rewards.
@@
$G_{t}=\sum_{j=t}^{T}\gamma^{j-t}r_{j}.$
Each RV faces a binary decision at every time step: Stop or Go. A Stop action holds the vehicle at the intersection entrance, while Go allows it to proceed. The RV makes this decision based on an observation vector $o_{t}$, which encodes traffic information across all directions:
$o_{t}=\bigoplus_{d\in D}\langle q_{d},\tau_{d}\rangle\bigoplus_{d\in D}\sigma_{d},$
@@
$r_{t}=\beta\cdot r^{\text{local}}_{t}+r^{\text{penalty}}_{t},$
where $\beta$ adjusts the importance of local efficiency. The local reward depends on the action: a Go yields a positive reward equal to $\tau_{d}$, while Stop incurs a negative reward of $-\tau_{d}$, representing delay. Additionally, if a conflict arises (i.e., multiple RVs attempt to enter the intersection simultaneously from conflicting directions), the vehicle receives a penalty of $-1$; otherwise, no penalty is applied.
@@ IV-A Experiment set-up
Our experimental setup consists of two unsignalized intersections managed by RVs and 12 signalized intersections controlled by traffic lights. For training, we apply the Rainbow DQN algorithm [36], conducting 1,000 episodes per configuration. RV penetration rates are varied at 25%, 50%, 75%, and 100%. Training times range from 8 to 16 hours per configuration on a machine equipped with an Intel Core™ i9-14900K CPU and an NVIDIA RTX 4080 GPU. We use a prioritized replay buffer with $\alpha=0.5$ and a capacity of 50,000. The neural network architecture consists of three hidden layers, each with 512 neurons, and employs 51 atoms for distributional output. Key hyperparameters include a discount factor of 0.99, a mini-batch size of 32, a learning rate of 0.0005, and a control zone radius of 30 meters.
Each trained policy is evaluated over 100 simulation runs, each lasting 1,000 seconds. Final performance metrics are computed by averaging results across these evaluations. During training, the network operates with 8,000 vehicles.
@@
    Experiments 7 vs. 8: Investigate how distinct OD flow patterns affect the overall average waiting time.
@@ IV-B Results and Analysis
Overall, the experimental results in Fig. 1 and Tab. I show that a uniform distribution of traffic across all 12 OD pairs significantly lowers average waiting time, outperforming all other configurations.
@@
These results highlight the importance of both traffic flow distribution and directionality in managing congestion. Notably, adjusting RV penetration alone does not guarantee consistent improvements across all scenarios. While higher RV rates generally help, the optimal performance is achieved when traffic is evenly distributed, especially at 75% RV penetration- average waiting time of 2.22 $s$. In contrast, the NE + SW configuration in 25% RV penetration yields the worst result, with an average waiting time of 18.11 $s$. In general, a uniform OD flow distribution has been shown to be most effective in minimizing congestion.
@@ V Conclusion and Future Work
We apply MARL to large-scale mixed traffic intersection control, conducting eight quasi-realistic experiments on a real-world urban network in Colorado Springs, USA, comprising 14 intersections. By exploring different OD flow patterns, we evaluate their influence on network-wide traffic performance. Results show that a uniform distribution of traffic across all 12 OD directions consistently yields the lowest average waiting time. Additionally, several OD configurations with directional emphasis also lead to improved performance, highlighting the potential of OD pattern design as an effective strategy for mitigating urban traffic congestion.
This work has several limitations. We only consider a fixed configuration of two unsignalized and twelve signalized intersections, without evaluating alternative distributions (e.g., 6 unsignalized + 8 signalized). We also focus exclusively on standard intersection geometries, excluding more complex layouts such as roundabouts and highway ramps. Furthermore, the influence of RV penetration rates shows non-linear patterns that merit deeper analysis. Future work will address these limitations by incorporating a wider range of intersection types and geometries, and by systematically analyzing the effects of varying RV penetration rates. These extensions aim to improve the robustness and generalizability of MARL-based traffic control strategies for real-world urban environments. There exist many future directions. At the macroscopic level, we can integrate this work with city-wide traffic reconstruction and prediction studies [37, 38, 39, 40, 41, 42, 43, 44, 45]; at the microscopic level, we can incorporate advanced robot vehicle motion planning and simulation methods [46, 47, 48, 49, 50, 51] into mixed traffic. The efforts at these two levels can potentially enable numerous Intelligent Transportation Systems applications.
[…]
Experiment RV rate
100% 75% 50% 25% Baseline
$\overline{W}(s)$
1 NS+SN 70% 14.69 16.00 15.09 7.83 11.80
2 NS+SN 90% 6.69 7.89 11.35 8.56 7.97
3 NW+WN 70% 14.60 17.80 12.48 13.46 11.12
4 NW+WN 90% 12.18 9.35 9.21 10.34 8.11
5 12 Directions in Uniform Distribution 5.09 2.22 2.66 4.71 2.36
6 WE+EW 70% 16.35 15.43 16.50 14.96 12.66
7 NE+SW 70% 16.81 15.62 14.17 18.11 12.51
8 SE+NW 70% 13.87 13.96 12.38 13.31 10.28
Table I: Average waiting time $\overline{W}$ of the eight experiments from 1) to 8). The experimental results indicate that a uniform distribution of traffic across all 12 OD directions significantly reduces average waiting times within the traffic network compared to other configurations. Specifically, when the primary traffic flow follows straight-line OD patterns such as NS+SN, increasing the emphasis on the direction from 70% to 90% leads to a decrease in average waiting time. Similarly, for a non-straight-line OD pattern like NW+WN, a higher traffic concentration (from 70% to 90%) in this direction also reduces waiting time. When comparing non-straight-line OD patterns (e.g., NW+WN) to straight-line patterns (e.g., NS+SN), there is no major difference. Among straight-line OD configurations, the NS+SN pattern contributes better performance than the WE+EW pattern overall. When evaluating different combinations of OD patterns, the SE+NW pattern demonstrates better outcomes than the SW+NE pattern.

### v2

@@ Abstract
Traffic congestion remains a major challenge for modern urban transportation, diminishing both efficiency and quality of life. While autonomous driving technologies and reinforcement learning (RL) have shown promise for improving traffic control, most prior work has focused on small-scale networks or isolated intersections. Large-scale mixed traffic control, involving both human-driven and robotic vehicles, remains underexplored. In this study, we propose a decentralized multi-agent reinforcement learning framework for managing large-scale mixed traffic networks, where intersections are controlled either by traditional traffic signals or by robotic vehicles. We evaluate our approach on a real-world network of 14 intersections in Colorado Springs, Colorado, USA, using average vehicle waiting time as the primary measure of traffic efficiency. We are exploring a problem that has not been sufficiently addressed: Is large-scale Multi-Agent Traffic Control (MTC) still feasible when facing time-varying Origin-Destination (OD) patterns?
@@ I Introduction
Traffic control has advanced considerably in recent years, with reinforcement learning (RL) emerging as a powerful tool in complex traffic scenarios. RL-based vehicle control strategies often outperform traditional traffic signal methods and have recently been extended to large-scale mixed traffic environments, especially for urban intersection management [5, 6]. Among these, multi-agent reinforcement learning (MARL) stands out for its strong performance and novel approach, attracting growing research interest [7, 8].
We conduct experiments on a large-scale traffic network of 14 intersections using the Simulation of Urban MObility (SUMO)[9]. In this network, some intersections are managed by traffic signals, while others are controlled by RVs equipped with RL capabilities. Our performance evaluation metric is average vehicle waiting time, which measures traffic efficiency by averaging the total time vehicles remain motionless near an intersection. We conduct experiments under various OD flow configurations and RV penetration rates. The results indicate that variations in OD configurations, when combined with appropriate RV penetration rates, can lead to significant differences in congestion levels. These findings suggest that adjusting the primary OD flow directions could serve as a viable strategy for alleviating urban traffic congestion. Such an approach opens up new avenues for dynamic traffic management and offers valuable insights for the future development of intelligent transportation systems. Results demonstrate that strategically adjusting major OD flow patterns can effectively influence congestion, offering a new pathway for enhancing urban mobility. This indicates that large-scale Multi-Agent Traffic Control (MTC) remains highly feasible even under time-varying OD patterns. Our work fills a gap in the current research and provides a more practical perspective for the future implementation of MTC in modern transportation systems in the real world. The code of our work is available at https://github.com/cgchrfchscyrh/MixedTrafficControl_IROS.
@@ II Related Work
Previous studies on unsignalized intersection control have focused primarily on connected and autonomous vehicles (CAVs). Early work introduced a multi-agent system in which CAVs reserve space-time slots on a first-come, first-served (FCFS) basis [14]. Later research expanded FCFS to CAV platoons [15] and proposed dynamic gap adjustments to reduce conflicts [16]. Decentralized approaches, like consensus-based trajectory planning [17] and energy optimization [18], have also been explored, but most assume a fully autonomous environment. In contrast, we address mixed traffic by assuming only RVs can communicate across intersections, offering a more realistic and scalable control strategy for real-world urban networks.
@@ III-A Mixed Traffic Control at Intersections Based on RL
$\mathcal{M}=\left(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O},\mathcal{Z},\mu_{0},\gamma\right)$
@@
  • $\mathcal{S}$: State space (set of all possible traffic states),
  • $\mathcal{A}$: Action space (e.g., Go or Stop),
  • $\mathcal{T}(s^{\prime}|s,a)$: Transition probability function,
  • $\mathcal{R}(s,a)$: Reward function,
  • $\mathcal{O}$: Observation space (for robot vehicles),
  • $\mathcal{Z}(o|s)$: Observation probability,
  • $\mu_{0}$: Initial state distribution,
  • $\gamma\in[0,1]$: Discount factor.
@@
$G_{t}=\sum_{j=t}^{T}\gamma^{j-t}r_{j}.$ (1)
Our action space is binary. Each RV faces a binary decision at every time step: Stop or Go. A Stop action holds the vehicle at the intersection entrance, while Go allows it to proceed. The RV makes this decision based on an observation vector $o_{t}$, which encodes traffic information across all directions:
$o_{t}=\bigoplus_{d\in D}\langle q_{d},\tau_{d}\rangle\bigoplus_{d\in D}\sigma_{d},$ (2)
@@
$r_{t}=\beta\cdot r_{t}^{\text{local}}+r_{t}^{\text{penalty}}$ (3)
The local reward is defined based on the action space discussed above :
$r_{t}^{\text{local}}=\begin{cases}+\tau_{d}&\text{if action is }\text{Go}\\ -\tau_{d}&\text{if action is }\text{Stop}\end{cases}$
The penalty term accounts for conflicts:
$r_{t}^{\text{penalty}}=\begin{cases}-1&\text{if a conflict occurs}\\ 0&\text{otherwise}\end{cases}$
@@ IV-A Experiment set-up
Our experimental setup consists of two unsignalized intersections managed by RVs and 12 signalized intersections controlled by traffic lights. For training, we apply the Rainbow DQN algorithm [29], conducting 1,000 episodes per configuration. Rainbow DQN is a reinforcement learning algorithm that integrates several improvements over the basic Deep Q-Network (DQN). It extends the original DQN by incorporating advancements such as Double DQN, Dueling Network Architecture, and Distributional RL. These enhancements collectively improve the algorithm’s ability to provide adaptive policies for heterogeneous agents, optimal decision-making, efficient learning under sparse rewards, and robustness in uncertain scenarios.  [29] In this work, we adopt a simplified yet efficient version of Rainbow DQN and apply it within a multi-agent setting, where all robotic vehicles (RVs) share a single policy network but make decisions independently. To implement Rainbow DQN, we use the standard DQN configuration with dueling and double Q-learning mechanisms enabled, and we conduct training within a multi-agent Ray framework. Additionally, we define the objective function for both the policy and the Q-function to support the learning process of Rainbow DQN as follows.
For the objective function defined below, the goal of each agent is to maximize the expected cumulative discounted reward over time:
$\max_{\pi_{\theta}}\mathbb{E}\left[\sum_{j=t}^{T}\gamma^{j-t}r_{j}\right]$ (4)
where:
  • $\pi_{\theta}(a_{t}|o_{t})$ is the shared policy of all RVs, parameterized by $\theta$.
  • $r_{j}$ is the reward received at time step $j$.
  • $\gamma$ is the discount factor (set to $0.99$ in this paper).
  • $T$ is the episode horizon (1000 seconds).
For the Q-function it is defined as:
$Q^{\pi}(o_{t},a_{t})=\mathbb{E}_{\pi}\left[R_{t}\mid o_{t},a_{t}\right],\quad\text{where}\quad R_{t}=\sum_{k=0}^{T-t}\gamma^{k}r_{t+k}$ (5)
Here:
  • $o_{t}$ denotes the local observation received by the agent at time step $t$;
  • $a_{t}$ is the action selected at time $t$, where $a_{t}\in\{\texttt{Go},\texttt{Stop}\}$;
  • $r_{t+k}$ is the reward obtained at future step $t+k$;
  • $\gamma\in(0,1)$ is the discount factor applied to future rewards (set to $\gamma=0.99$);
  • $\pi$ represents the shared stochastic policy followed by all agents after time $t$.
Because the environment is modeled as a decentralized partially observable Markov decision process (Dec-POMDP), each agent estimates its Q-value based solely on partial local information, without access to the full environment state. To learn the Q-function, the Rainbow DQN algorithm is used, which incorporates several state-of-the-art DQN improvements:
  • Double Q-learning: mitigates overestimation bias by separating the action selection (from the online network) and value evaluation (using the target network).
  • Dueling Network Architecture: decomposes the Q-value function into two separate estimators:
    $Q(o,a)=V(o)+\left(A(o,a)-\frac{1}{|\mathcal{A}|}\sum_{a^{\prime}\in\mathcal{A}}A(o,a^{\prime})\right)$ (6)
    where $V(o)$ is the estimated state value, and $A(o,a)$ is the advantage of taking action $a$ under observation $o$.
  • Distributional RL: Instead of estimating a scalar Q-value, it learns the full probability distribution of possible returns using 51 discrete support atoms.
RV penetration rates are varied at 25%, 50%, 75%, and 100%. Training times range from 8 to 16 hours per configuration on a machine equipped with an Intel Core™ i9-14900K CPU and an NVIDIA RTX 4080 GPU. We use a prioritized replay buffer with $\alpha=0.5$ and a capacity of 50,000. The neural network architecture consists of three hidden layers, each with 512 neurons, and employs 51 atoms for distributional output. Key hyperparameters include a discount factor of 0.99, a mini-batch size of 32, a learning rate of 0.0005, and a control zone radius of 30 meters. Each trained policy is evaluated over 100 simulation runs, each lasting 1,000 seconds, with 8,000 vehicles running in the network. Final performance metrics are computed by averaging results across these evaluations.
@@
    Experiments 7 vs. 8: Investigate how different OD flow patterns affect the overall average waiting time.
@@ IV-B Results and Analysis
The experimental results in Fig. 1 and Table I show that a uniform traffic distribution across all 12 OD pairs significantly reduces average waiting time, outperforming all other configurations.
RV rate
100% 75% 50% 25% Baseline
Experiment $\overline{W}(s)$
1. NS+SN 70% 14.69 16 15.09 7.83 11.8
2. NS+SN 90% 6.69 7.89 11.35 8.56 7.97
3. NW+WN 70% 14.6 17.8 12.48 13.46 11.12
4. NW+WN 90% 12.18 9.35 9.21 10.34 8.11
5. 12 Dir. Uniform. 5.09 2.22 2.66 4.71 2.36
6. WE+EW 70% 16.35 15.43 16.50 14.96 12.66
7. NE+SW 70% 16.81 15.62 14.17 18.11 12.51
8. SE+NW 70% 13.87 13.96 12.38 13.31 10.28
Table I: Average waiting time $\overline{W}$ of the eight experiments from 1) to 8). “12 Dir Uniform” means 12 directions in uniform distribution. The experimental results indicate that a uniform distribution of traffic across all 12 OD directions significantly reduces average waiting times within the traffic network compared to other configurations. Specifically, when the primary traffic flow follows straight-line OD patterns such as NS+SN, increasing the emphasis on the direction from 70% to 90% leads to a decrease in average waiting time. Similarly, for a non-straight-line OD pattern like NW+WN, a higher traffic concentration (from 70% to 90%) in this direction also reduces waiting time. When comparing non-straight-line OD patterns (e.g., NW+WN) to straight-line patterns (e.g., NS+SN), there is no major difference. Among straight-line OD configurations, the NS+SN pattern contributes better performance than the WE+EW pattern overall. When evaluating different combinations of OD patterns, the SE+NW pattern demonstrates better outcomes than the SW+NE pattern.
@@
These results highlight the importance of both traffic flow distribution and directionality in managing congestion. Notably, adjusting RV penetration alone does not guarantee consistent improvements across all scenarios. While higher RV rates generally help, the optimal performance is achieved when traffic is evenly distributed, especially at 75% RV penetration- average waiting time of 2.22 $s$. In contrast, the NE + SW configuration in 25% RV penetration yields the worst result, with an average waiting time of 18.11 $s$. A uniform OD flow distribution is most effective in minimizing congestion.
@@ V Conclusion and Future Work
We apply MARL to large-scale mixed traffic intersection control, conducting eight quasi-realistic experiments on a real-world urban network in Colorado Springs, USA, comprising 14 intersections. By exploring different OD flow patterns, we evaluate their influence on network-wide traffic performance. Results show that a uniform distribution of traffic across all 12 OD directions consistently yields the lowest average waiting time. Additionally, several OD configurations with directional emphasis also lead to improved performance, highlighting the potential of OD pattern design as an effective strategy for mitigating urban traffic congestion and proving that large-scale Multi-Agent Traffic Control (MTC) still holds potential when facing time-varying OD patterns.
This work has several limitations. We only consider a fixed configuration of two unsignalized and twelve signalized intersections, without evaluating alternative distributions (e.g., 6 unsignalized + 8 signalized). We also focus exclusively on standard intersection geometries, excluding more complex layouts such as roundabouts and highway ramps. Furthermore, the influence of RV penetration rates shows non-linear patterns that merit deeper analysis. Future work will address these limitations by incorporating a wider range of intersection types and geometries, and by systematically analyzing the effects of varying RV penetration rates. These extensions aim to improve the robustness and generalizability of MARL-based traffic control strategies for real-world urban environments. The transition from intersections fully controlled by traffic signals to a system without signals is a gradual process. Our research can serve as a theoretical foundation for pilot testing of signal-free intersections, starting with a single intersection that operates without traffic lights.
Acknowledgment
This research is funded by the National Science Foundation (NSF) via Grant 2129003, 2231710, and 2153426. The authors gratefully acknowledge NSF’s support. The authors would also like to thank NVIDIA and the Tickle College of Engineering at University of Tennessee, Knoxville for their support.